5. a) Calculate the area under the curve y = x ^ 3 + 4x + 1 from x = - 3 to x = 3 .


b) Evaluate: integrate x * e ^ x dx from 0 to 4

1
Expert's answer
2021-12-09T10:08:18-0500

(a) "y = x ^ 3 + 4x + 1"



Graph will intersect the curve at "x=-0.25"

Area under the curve (green region) = "\int_{-0.25}^{3}(x^3+4x+1)dx-\int_{-3}^{-0.25}(x^3+4x+1)dx"

"(x^4/4+4x^2/2+x)_{-0.25}^{3} \ - (x^4/4+4x^2/2+x)_{-3}^{-0.25} \\
=27.89+21.13+3.25-(14.29+15.13+2.75)\\
=20.1"


(b) "\int_{0}^{4} xe^xdx= [xe^x-\int (e^x.1)dx]_{0}^{4}\\"

"=[xe^x-\int (e^x.1)dx]_{0}^{4}=[xe^x-e^x]_{0}^{4}=4.e^4-e^4-0+1=54.59"